# Learning in complex action spaces without policy gradients

### v1

@@ Abstract
Conventional wisdom suggests that policy gradient methods are better suited to complex action spaces than action-value methods. However, foundational studies have shown equivalences between these paradigms in small and finite action spaces (O’Donoghue et al., 2017; Schulman et al., 2017a). This raises the question of why their computational applicability and performance diverge as the complexity of the action space increases. We hypothesize that the apparent superiority of policy gradients in such settings stems not from intrinsic qualities of the paradigm, but from universal principles that can also be applied to action-value methods to serve similar functionality. We identify three such principles and provide a framework for incorporating them into action-value methods. To support our hypothesis, we instantiate this framework in what we term QMLE, for Q-learning with maximum likelihood estimation. Our results show that QMLE can be applied to complex action spaces with a controllable computational cost that is comparable to that of policy gradient methods, all without using policy gradients. Furthermore, QMLE demonstrates strong performance on the DeepMind Control Suite, even when compared to the state-of-the-art methods such as DMPO and D4PG.
@@ 1 Introduction
In reinforcement learning, policy gradients have become the backbone of solutions for environments with complex action spaces, including those involving large, continuous, combinatorial, or structured subaction spaces (Dulac-Arnold et al., 2015; OpenAI et al., 2019; Vinyals et al., 2019; Hubert et al., 2021; Ouyang et al., 2022). In contrast, action-value methods have traditionally been confined to tabular-action models for small and finite action spaces. However, where applicable, such as on the Atari Suite (Bellemare et al., 2013; Machado et al., 2018), action-value methods are frequently the preferred approach over policy gradient methods (Kapturowski et al., 2023; Schwarzer et al., 2023).
@@
Over the past years, foundational research has shown that the distinction between action-value and policy gradient methods is narrower than previously understood, particularly in the basic case of tabular-action models in small and finite action spaces (see, e.g., Schulman et al., 2017a). Notably, O’Donoghue et al. (2017) established an equivalency between these paradigms, revealing a direct connection between the fixed-points of the action-preferences of policies optimized by regularized policy gradients and the action-values learned by action-value methods. These insights invite further exploration of the discrepancies that emerge as the complexity of action spaces increases.
@@
What are the core principles that underpin the greater computational applicability and performance of policy gradient methods in such settings? In this paper, we identify three such principles. First, policy gradient methods leverage Monte Carlo (MC) approximations for summation or integration over the action space, enabling computational feasibility even in environments with complex action spaces. Second, they employ amortized maximization through a special form of maximum likelihood estimation (namely, the policy gradient itself), iteratively refining the policy to increase the likelihood of selecting high-value actions without requiring brute-force $\operatorname*{arg\,max}$ over the action space. Third, scalable policy gradient methods employ action-in architectures for action-value approximation, which covertly enable representation learning and generalization across the joint state-action space.
@@
Are these principles exclusive to policy gradient methods? We argue that these principles can be adapted to action-value methods. Specifically, instead of using MC methods for summation or integration as in policy gradient methods, they can be used to approximate the $\operatorname*{arg\,max}$ in action-value methods in order to make them computationally scalable for complex action spaces. Moreover, explicit maximum likelihood estimation can be applied to enable caching and iterative refinement of parametric predictors for amortized $\operatorname*{arg\,max}$ approximation. Lastly, action-in architectures can be employed not only as a scalable approach for evaluating a limited set of actions in any given state, but also to enable representation learning and generalization across both states and actions.
We introduce $Q$-learning with maximum likelihood estimation (QMLE) to test our hypotheses. Our empirical study shows that QMLE achieves strong performance in environments with complex action spaces, all while matching the computational complexity of policy gradient methods. These results provide evidence that the identified principles are core to the success of policy gradient methods in such environments. Moreover, they support that the principles are not intrinsic to the policy gradient paradigm, but are universal and adaptable to action-value learning for achieving similar qualities.
@@ 2.1 The reinforcement learning problem
The goal of an RL agent is defined as finding an optimal policy $\pi^{*}$ that maximizes this quantity across the state space: $V^{\pi^{*}}\geq V^{\pi}$ for all $\pi$. While there may be more than one optimal policy, they all share the same state-value function: $V^{*}=V^{\pi^{*}}$. Similarly, we can define the value of state $s$ and action $a$ under policy $\pi$: $Q^{\pi}(s,a)\doteq\mathbb{E}_{\pi,\mathcal{P},\mathcal{R}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}|s_{0}=s,a_{0}=a]$. Notice that the goal can be equivalently phrased as finding an optimal policy $\pi^{*}$ that maximizes this alternative quantity across the joint state-action space: $Q^{\pi^{*}}\geq Q^{\pi}$ for all $\pi$. Same as before, optimal policies share the same action-value function: $Q^{*}=Q^{\pi^{*}}$.
The state and action value functions are related to each other via: $V^{\pi}(s)=\sum_{a}Q^{\pi}(s,a)\pi(a|s)$, where we use $\sum$ to signify both summation and integration over discrete or continuous actions. For all MDPs there is always at least one deterministic optimal policy, which can be deduced by maximizing the optimal action-value function: $\operatorname*{arg\,max}_{a}Q^{*}(s,a)$ in any given state $s$. It is worth noting that there may be cases where multiple actions yield the same maximum value, resulting in ties. By breaking such ties at random, considering all conceivable distributions, we can construct the set of all optimal policies, including both deterministic and stochastic policies. Regardless of the optimal policy, the optimal state-value and action-value functions are related to each other in the following way: $V^{*}(s)=\max_{a}Q^{*}(s,a)$. Similarly, the optimal state-value function can be used to extract optimal policies by invoking the Bellman recurrence: $\operatorname*{arg\,max}_{a}\mathbb{E}_{\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma V^{*}(s_{t+1})|s_{t}=s]$. However, this requires access to the MDP model, rendering the sole optimization of state-values unsuitable for model-free RL.
@@ 2.2 Action-value learning
$Q^{\pi}(s,a)=\mathbb{E}_{\pi,\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma Q^{\pi}(s_{t+1},a_{t+1})|s_{t}=s,a_{t}=a],$ (1)
@@
$Q^{*}(s,a)=\mathbb{E}_{\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma\max_{a^{\prime}}Q^{*}(s_{t+1},a^{\prime})|s_{t}=s,a_{t}=a].$ (2)
@@ 2.3 Policy gradient methods
$\pi^{*}\doteq\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}.$ (5)
@@
$\nabla\mathbb{E}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}=\nabla\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\pi(a|s_{t})\Big{]}\propto\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\nabla\pi(a|s_{t})\Big{]}.$ (6)
@@
$\nabla\mathbb{E}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}$ $=\nabla\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\int Q^{\pi}(s_{t},a)\delta\big{(}a-\pi(s_{t})\big{)}\textrm{d}a\Big{]}$ (7a)
$=\nabla\mathbb{E}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}\big{(}s_{t},\pi(s_{t})\big{)}\Big{]}$ (7b)
$\propto\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\nabla_{a}Q^{\pi}\big{(}s_{t},a\!=\!\pi(s_{t})\big{)}\nabla\pi(s_{t})\Big{]}.$ (7c)
@@
$\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\nabla_{a}Q^{\mu}\big{(}s_{t},a\!=\!\mu(s_{t})\big{)}\nabla\mu(s_{t})\Big{]},$ (8)
@@ 2.4 Maximum likelihood estimation
Suppose we have a data set $\{(x_{i},y_{i})\}$ drawn from an unknown joint distribution $p(x,y)$, where random variables $x_{i}$ and $y_{i}$ respectively represent inputs and targets. Frequently, problem scenarios involve determining the parameters of an assumed probability distribution that best describe the data. The method of maximum likelihood estimation (MLE) addresses this by posing the question: “Under which parameter values is the observed data most likely?”. In this context, we typically start by representing our assumed distribution using a parameterized function $f$, such as a neural network, with parameters $\bm{\mathbf{\theta}}$. Hence, $\bm{\mathbf{\phi}}\doteq f(x)$ serves as our estimator for the distributional parameters in $x$. For example, $\bm{\mathbf{\phi}}$ contains $K$ values in the case of a categorical distribution with $K$ categories, and contains means $\bm{\mathbf{\mu}}$ and variances $\bm{\mathbf{\sigma}}$ in the case of a multivariate heteroscedastic Gaussian distribution. We will denote the probability distribution that is specified by parameters $\bm{\mathbf{\phi}}=f(x)$ as $f(y|x)$. The problem of finding the optimal parameters can then be formulated as:
$\operatorname*{arg\,max}_{\bm{\mathbf{\phi}}}\mathbb{E}_{p(x,y)}\Big{[}\log f(y_{i}|x_{i})\Big{]}.$ (10)
@@ 3.1 Approximate summation or integration using Monte Carlo methods
$\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\nabla\pi(a|s_{t})\Big{]}$ $=\mathbb{E}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}(s_{t},a_{t})\frac{\nabla\pi(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\Big{]}$ (12)
$=\mathbb{E}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}(s_{t},a_{t})\nabla\log\pi(a_{t}|s_{t})\Big{]},$
@@
$\frac{1}{n}\sum_{t}{\color[rgb]{0.33203125,0.0390625,0.78515625}\frac{1}{m+1}}\Big{(}Q^{\pi}(s_{t},a_{t})\nabla\log\pi(a_{t}|s_{t}){\color[rgb]{0.33203125,0.0390625,0.78515625}~{}+\sum_{i=0}^{m-1}Q^{\pi}(s_{t},a_{i})\nabla\log\pi(a_{i}|s_{t})}\Big{)},$ (14)
where $m$ is the number of off-trajectory action samples $a_{i}\sim\pi(.|s_{t})$ per state $s_{t}$. When $m=0$, this reduces to the original on-trajectory policy gradient estimator (13). It is important to note that using the policy gradient estimator (14) with $m>0$ requires direct approximation of the action-values $Q^{\pi}$ by a function $Q$, e.g. a neural network trained by minimizing the squared form of the TD residual (3).
@@ 3.2 Amortized maximization using maximum likelihood estimation
We start with recasting the log-likelihood gradient (11) using RL terminology, replacing the variables $(x,y,i,f)$ with $(s,a,t,\pi)$. Moreover, we reinterpret the expectation computation to be under the joint visitation distribution of state-action pairs within an RL context. Subsequently, we contrast the reframed log-likelihood gradient against the policy gradient (12):
$\underbrace{\mathbb{E}_{\pi,\mathcal{P}}\Big{[}\nabla\log\pi(a_{t}|s_{t})\Big{]}}_{\textrm{{{\color[rgb]{0.33203125,0.0390625,0.78515625}log-likelihood gradient}}}}\;\;\;\;\textrm{vs.}\;\;\;\;\underbrace{\mathbb{E}_{\pi,\mathcal{P}}\Big{[}{\color[rgb]{0.33203125,0.0390625,0.78515625}Q^{\pi}(s_{t},a_{t})}\nabla\log\pi(a_{t}|s_{t})\Big{]}}_{\textrm{{{\color[rgb]{0.33203125,0.0390625,0.78515625}policy gradient}}}}.$ (15)
We have discussed that policy gradients can be viewed as an iterative approach to action-value maximization. However, they do not always yield the global $\operatorname*{arg\,max}$. This limitation is rooted in local tendencies of gradient-based optimization, affecting scenarios with non-tabular policy distributions (Tessler et al., 2019). Figures 1c,d respectively show progression of a delta policy and a fixed-variance Gaussian policy in a continuous bandit problem with bimodal and deterministic rewards. In both cases, policy improvement driven by policy gradients results in local movement in the action space and thus convergence to sub-optimal policies.

### v2

@@ Abstract
While conventional wisdom holds that policy gradient methods are better suited to complex action spaces than action-value methods, foundational work has shown that the two paradigms are equivalent in small, finite action spaces (O’Donoghue et al., 2017; Schulman et al., 2017a). This raises the question of why their computational applicability and performance diverge as the complexity of the action space increases. We hypothesize that the apparent superiority of policy gradients in such settings stems not from intrinsic qualities of the paradigm but from universal principles that can also be applied to action-value methods, enabling similar functions. We identify three such principles and provide a framework for incorporating them into action-value methods. To support our hypothesis, we instantiate this framework in what we term QMLE, for Q-learning with maximum likelihood estimation. Our results show that QMLE can be applied to complex action spaces at a computational cost comparable to that of policy gradient methods, all without using policy gradients. Furthermore, QMLE exhibits strong performance on the DeepMind Control Suite, even when compared to state-of-the-art methods such as DMPO and D4PG. We make our code publicly available.
@@ 1 Introduction
In reinforcement learning, policy gradients (and their actor-critic variants) (Sutton et al., 1999) have become the backbone of solutions for environments with complex action spaces (Dulac-Arnold et al., 2015; OpenAI et al., 2019; Vinyals et al., 2019; Ouyang et al., 2022). We use the term “complex action spaces” broadly to refer to those for which enumerating actions or brute-force operations are computationally intractable, such as domains with multi-dimensional continuous or high-dimensional discrete actions (Hubert et al., 2021). In contrast, action-value methods—primarily, variants of Sarsa and Q-learning—have traditionally been confined to tabular-action models for small and finite action spaces. However, where applicable, such as on the Atari Suite (Bellemare et al., 2013; Machado et al., 2018), action-value methods are frequently the preferred approach over policy gradient methods (Kapturowski et al., 2023; Schwarzer et al., 2023).
@@
In recent years, foundational research has shown that the distinction between action-value and policy gradient methods is narrower than previously understood, particularly in the basic case of tabular-action models in small and finite action spaces (see, e.g., Schulman et al., 2017a). Notably, O’Donoghue et al. (2017) established a direct equivalence between these paradigms, connecting the fixed-points of action-preferences of policies optimized by regularized policy gradients to action-values learned by action-value methods. These insights invite further exploration of discrepancies emerging as action-space complexity increases.
@@
What core principles underpin the superior computational applicability and performance of policy gradient methods in such settings? In this paper, we identify three such principles. First, policy gradient methods leverage Monte Carlo (MC) approximations for summation or integration over action spaces, enabling computational feasibility even in environments with complex action spaces. Second, they employ amortized maximization through a special form of maximum likelihood estimation (namely, the policy gradient itself), iteratively refining the policy to favor high-value actions without requiring brute-force $\operatorname*{arg\,max}$ over the action space. Third, scalable policy gradient methods employ action-in architectures for action-value approximation, implicitly facilitating representation learning and generalization across the joint state-action space.
@@
Are these principles exclusive to policy gradient methods? We argue that these principles can be adapted to action-value methods, thereby bridging the scalability and performance gap between the two paradigms. Instead of using MC methods for summation or integration as in policy gradient methods, we propose using MC methods to approximate the $\operatorname*{arg\,max}$ operation, making action-value methods computationally scalable for complex action spaces. Moreover, explicit maximum likelihood estimation can enable caching and iterative refinement of parametric predictors for amortized $\operatorname*{arg\,max}$ approximation. Lastly, action-in architectures can be employed not only to scalably evaluate a limited set of actions in any given state, but also to enable representation learning and generalization across both states and actions.
To test our arguments, we introduce $Q$-learning with maximum likelihood estimation (QMLE). Given that QMLE builds upon Q-learning, DDPG (Lillicrap et al., 2016) serves as its closest counterpart among commonly used policy gradient methods for continuous-action tasks, and thus is our primary baseline for comparison.
Through systematic ablation studies (§5.3), we show that each identified principle individually contributes to improved scalability or performance of action-value learning. Benchmarking QMLE against DDPG on diverse multi-dimensional continuous control tasks, ranging from 1 to 21 action dimensions, we demonstrate that incorporating all three principles collectively enables QMLE to achieve competitive performance (§5.2). These results provide evidence that the identified principles are indeed fundamental to the broad applicability of policy gradient methods in complex action spaces, and that they are not intrinsically tied to the policy gradient paradigm and can be effectively adapted into action-value learning.
Beyond multi-dimensional continuous-action scenarios, we demonstrate the applicability of QMLE to large, high-dimensional discrete-action ones by employing a 3-bin discretization scheme on our continuous control benchmarks. Our results illustrate QMLE’s robustness in managing increasing complexity due to the curse of dimensionality, effectively scaling from 2 to 38 action dimensions (corresponding to discrete action sets ranging from only $3^{2}=9$ up to $3^{38}$, approximately 1.35 quintillion, actions).
@@ 2.1 The reinforcement learning problem
The goal of an RL agent is defined as finding an optimal policy $\pi^{*}$ that maximizes this quantity across the state space: $V^{\pi^{*}}\geq V^{\pi}$ for all $\pi$. While there may be more than one optimal policy, they all share the same state-value function: $V^{*}=V^{\pi^{*}}$. Similarly, we can define the value of state $s$ and action $a$ under policy $\pi$: $Q^{\pi}(s,a)\doteq\operatorname*{\mathbb{E}}_{\pi,\mathcal{P},\mathcal{R}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}|s_{0}=s,a_{0}=a]$. Notice that the goal can be equivalently phrased as finding an optimal policy $\pi^{*}$ that maximizes this alternative quantity across the joint state-action space: $Q^{\pi^{*}}\geq Q^{\pi}$ for all $\pi$. Same as before, optimal policies share the same action-value function: $Q^{*}=Q^{\pi^{*}}$.
The state and action value functions are related to each other via: $V^{\pi}(s)=\sum_{a}Q^{\pi}(s,a)\pi(a|s)$, where we use $\sum$ to signify both summation and integration over discrete or continuous actions. For all MDPs there is always at least one deterministic optimal policy, which can be deduced by maximizing the optimal action-value function: $\operatorname*{arg\,max}_{a}Q^{*}(s,a)$ in any given state $s$. It is worth noting that there may be cases where multiple actions yield the same maximum value, resulting in ties. By breaking such ties at random, considering all conceivable distributions, we can construct the set of all optimal policies, including both deterministic and stochastic policies. Regardless of the optimal policy, the optimal state-value and action-value functions are related to each other in the following way: $V^{*}(s)=\max_{a}Q^{*}(s,a)$. Similarly, the optimal state-value function can be used to extract optimal policies by invoking the Bellman recurrence: $\operatorname*{arg\,max}_{a}\operatorname*{\mathbb{E}}_{\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma V^{*}(s_{t+1})|s_{t}=s]$. However, this requires access to the MDP model, rendering the sole optimization of state-values unsuitable for model-free RL.
@@ 2.2 Action-value learning
$Q^{\pi}(s,a)=\operatorname*{\mathbb{E}}_{\pi,\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma Q^{\pi}(s_{t+1},a_{t+1})|s_{t}=s,a_{t}=a],$ (1)
@@
$Q^{*}(s,a)=\operatorname*{\mathbb{E}}_{\mathcal{P},\mathcal{R}}[r_{t+1}+\gamma\max_{a^{\prime}}Q^{*}(s_{t+1},a^{\prime})|s_{t}=s,a_{t}=a].$ (2)
@@ 2.3 Policy gradient methods
$\pi^{*}\doteq\operatorname*{arg\,max}_{\pi}\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}.$ (5)
@@
$\nabla\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}=\nabla\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\pi(a|s_{t})\Big{]}\propto\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\nabla\pi(a|s_{t})\Big{]}.$ (6)
@@
$\nabla\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}V^{\pi}(s_{t})\Big{]}$ $=\nabla\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\int Q^{\pi}(s_{t},a)\delta\big{(}a-\pi(s_{t})\big{)}\textrm{d}a\Big{]}$ (7a)
$=\nabla\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}\big{(}s_{t},\pi(s_{t})\big{)}\Big{]}$ (7b)
$\propto\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\nabla_{a}Q^{\pi}\big{(}s_{t},a\!=\!\pi(s_{t})\big{)}\nabla\pi(s_{t})\Big{]}.$ (7c)
@@
$\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\nabla_{a}Q^{\mu}\big{(}s_{t},a\!=\!\mu(s_{t})\big{)}\nabla\mu(s_{t})\Big{]},$ (8)
@@ 2.4 Maximum likelihood estimation
Suppose we have a data set $\{(x_{i},y_{i})\}$ drawn from an unknown joint distribution $p(x,y)$, where random variables $x_{i}$ and $y_{i}$ respectively represent inputs and targets. Frequently, problem scenarios involve determining the parameters of an assumed probability distribution that best describe the data. The method of maximum likelihood estimation (MLE) addresses this by posing the question: “For which parameter values is the observed data most likely?”. In this context, we typically start by representing our assumed distribution using a parameterized function $f$, such as a neural network, with parameters $\bm{\mathbf{\theta}}$. Hence, $\bm{\mathbf{\phi}}\doteq f(x)$ serves as our estimator for the distributional parameters in $x$. For example, $\bm{\mathbf{\phi}}$ contains $K$ values in the case of a categorical distribution with $K$ categories, and contains means $\bm{\mathbf{\mu}}$ and variances $\bm{\mathbf{\sigma}}$ in the case of a multivariate heteroscedastic Gaussian distribution. We will denote the probability distribution that is specified by parameters $\bm{\mathbf{\phi}}=f(x)$ as $f(y|x)$. The problem of finding the optimal parameters can then be formulated as:
$\operatorname*{arg\,max}_{\bm{\mathbf{\phi}}}\operatorname*{\mathbb{E}}_{p(x,y)}\Big{[}\log f(y_{i}|x_{i})\Big{]}.$ (10)
@@ 3.1 Approximate summation or integration using Monte Carlo methods
$\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\sum_{a}Q^{\pi}(s_{t},a)\nabla\pi(a|s_{t})\Big{]}$ $=\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}(s_{t},a_{t})\frac{\nabla\pi(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\Big{]}$ (12)
$=\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}Q^{\pi}(s_{t},a_{t})\nabla\log\pi(a_{t}|s_{t})\Big{]},$
@@
$\frac{1}{n}\sum_{t}{\color[rgb]{0.33203125,0.0390625,0.78515625}\frac{1}{m+1}}\Big{(}Q^{\pi}(s_{t},a_{t})\nabla\log\pi(a_{t}|s_{t}){\color[rgb]{0.33203125,0.0390625,0.78515625}~+}{\color[rgb]{0.33203125,0.0390625,0.78515625}\mathbf{1}_{\{m>0\}}\sum_{i=0}^{m-1}Q^{\pi}(s_{t},a_{i})\nabla\log\pi(a_{i}|s_{t})}\Big{)},$ (14)
where $m$ is the number of off-trajectory action samples $a_{i}\sim\pi(.|s_{t})$ per state $s_{t}$, and the indicator term $\mathbf{1}_{\{m>0\}}$ ensures that when $m=0$, the second term vanishes, thus reducing to the original on-trajectory policy gradient estimator (13). It is important to note that using the on-trajectory estimator (13) applies broadly to both critic-free (e.g. REINFORCE) and actor-critic (e.g. PPO) methods, but using the off-trajectory estimator (14) with $m>0$ requires direct approximation of the action-values $Q^{\pi}$ by a function $Q$, e.g. a neural network trained by minimizing the squared form of the TD residual (3). As such, the latter can only be employed by $Q$-based actor-critic methods (e.g. DDPG).
@@ 3.2 Amortized maximization using maximum likelihood estimation
We start with recasting the log-likelihood gradient (11) using RL terminology, replacing the variables $(x,y,i,f)$ with $(s,a,t,\pi)$. Moreover, we reinterpret the expectation computation to be under the joint visitation distribution of state-action pairs within an RL context. [^5] While the reframed log-likelihood gradient is useful for comparison against the policy gradient, it does not per se specify a meaningful optimization problem. This is because the target distribution and its estimator are equivalent (i.e. $\pi(a|s)p(s)$). Subsequently, we contrast the reframed log-likelihood gradient against the policy gradient (12):
$\underbrace{\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}\nabla\log\pi(a_{t}|s_{t})\Big{]}}_{\textrm{\normalsize{{\color[rgb]{0.33203125,0.0390625,0.78515625}log-likelihood gradient}}}}\;\;\;\;\textrm{vs.}\;\;\;\;\underbrace{\operatorname*{\mathbb{E}}_{\pi,\mathcal{P}}\Big{[}{\color[rgb]{0.33203125,0.0390625,0.78515625}Q^{\pi}(s_{t},a_{t})}\nabla\log\pi(a_{t}|s_{t})\Big{]}}_{\textrm{\normalsize{{\color[rgb]{0.33203125,0.0390625,0.78515625}policy gradient}}}}.$ (15)
@@
We have discussed that policy gradients can be viewed as an iterative approach to action-value maximization. However, they do not always yield the global $\operatorname*{arg\,max}$. This limitation is rooted in local tendencies of gradient-based optimization, affecting scenarios with non-tabular policy distributions (Tessler et al., 2019). Figures 1c,d respectively show progression of a delta policy and a fixed-variance Gaussian policy in a continuous bandit with bimodal, deterministic rewards. In both cases, policy improvement driven by policy gradients results in local movement in the action space and thus convergence to suboptimal policies.